# Sufficient conditions for strong discrete maximum principles in finite element solutions of linear and semilinear elliptic equations

### v1

@@ 6. The semilinear problem
Consider the discrete semilinear elliptic equation under the formulation (2.13), and let $\tilde{c}\geq 0$ satisfy (2.3)–(2.4). Furthermore, assume there is a constant $C_{A}>0$ independent of $h$ so that whenever vertices $P_{i}$ and $P_{j}$ are adjacent with at least one of them lying in the interior (that is, $\overline{P_{i}P_{j}}$ is not a boundary edge), we have ${\mathbf{A}}_{ij}<0$ and
@@
We argument follows the same path as in Theorem 4.2, namely we prove that ${\mathcal{S}}^{E}_{h}$ satisfies the sDMP-B for any star $E_{P}$ around a vertex $P$. The conclusion will follow from Corollary 3.5.
@@
Indeed, given an interior edge $\overline{P_{i}P_{j}}$ in the notation from Section 4.1 we have cf. (4.5)
$\int_{{\mathcal{T}}}\varphi_{i}\varphi_{j}=\frac{h_{k}^{2}}{24(\cot\theta_{i}+\cot\theta_{j})}\leq\frac{h_{k}^{2}}{48\cot\theta_{0}}=\frac{h_{k}^{2}\tan\theta_{0}}{48},$
where $h_{k}=|\overline{P_{i}P_{j}}|$. Since the contribution to ${\mathbf{M}}_{ij}$ comes from two triangles, we have
${\mathbf{M}}_{ij}\leq\frac{h_{k}^{2}\tan\theta_{0}}{24}.$
@@
$\int_{{\mathcal{T}}_{1}\cup{\mathcal{T}}_{2}}\nabla\varphi_{i}\cdot\nabla\varphi_{i}=\frac{\sin(\theta_{1}+\theta_{2})}{2\sin\theta_{1}\sin\theta_{2}}=\frac{1}{2}(\cot\theta_{1}+\cot\theta_{2})\geq\cot\left(\frac{\pi}{2}-\theta_{0}\right)=\tan\theta_{0}.$
Due to the contribution of two adjacent triangles, we have
${\mathbf{A}}_{ij}\geq 2\tan\theta_{0}.$
@@
${\mathbf{M}}_{ij}\leq-\frac{h_{k}^{2}}{48}{\mathbf{A}}_{ij}.$
Hence, $C_{A}=1/48$.
@@ 7. Conclusions
Note that ${\mathbf{T}}_{0}$ is singular. The computation has been validated numerically against stiffness matrix computations using standard finite element codes. ∎

### v2

@@ 6. The semilinear problem
Consider the discrete semilinear elliptic equation under the formulation (2.14), and let ${c}\geq 0$ satisfy (2.3)–(2.5). Furthermore, assume there is a constant $C_{A}>0$ independent of $h$ so that whenever vertices $P_{i}$ and $P_{j}$ are adjacent with at least one of them lying in the interior (that is, $\overline{P_{i}P_{j}}$ is not a boundary edge), we have ${\mathbf{A}}_{ij}<0$ and
@@
The argument follows the same path as in Theorem 4.2, namely we prove that ${\mathcal{S}}^{E}_{h}$ satisfies the sDMP-B for any star $E_{P}$ around a vertex $P$. The conclusion will follow from Corollary 3.5.
@@
Indeed, given an interior edge $\overline{P_{i}P_{j}}$ in the notation from Section 4.1 and referring to Fig. 1 (right) we have cf. (4.5)
$\int_{\Delta P_{i}P_{j}P_{k}}\varphi_{i}\varphi_{j}=\frac{h_{k}^{2}}{24(\cot\theta_{i}+\cot\theta_{j})}\leq\frac{h_{k}^{2}}{48\cot\left(\frac{\pi}{2}-\alpha_{0}\right)}=\frac{h_{k}^{2}}{48\tan\alpha_{0}},$
where $h_{k}=|\overline{P_{i}P_{j}}|$. Since two adjacent triangles contribute to ${\mathbf{M}}_{ij}$, we have
${\mathbf{M}}_{ij}\leq\frac{h_{k}^{2}}{24\tan\alpha_{0}}.$
@@
${\mathbf{A}}_{ij}=\frac{\sin(\theta_{1}+\theta_{2})}{2\sin\theta_{1}\sin\theta_{2}}=\frac{1}{2}(\cot\theta_{1}+\cot\theta_{2})\geq\cot\left(\frac{\pi}{2}-\alpha_{0}\right)=\tan\alpha_{0}.$
@@
${\mathbf{M}}_{ij}\leq\frac{h_{k}^{2}\tan\alpha_{0}}{24\tan^{2}\alpha_{0}}\leq\frac{h_{k}^{2}}{24\tan^{2}\alpha_{0}}\>|{\mathbf{A}}_{ij}|.$
Hence, $C_{A}=(24\tan^{2}\alpha_{0})^{-1}$.
@@ 7. Conclusions
Note that ${\mathbf{T}}_{0}$ is singular, and has rank 3, hence is rank-3 deficient. The computation has been validated numerically against stiffness matrix computations using standard finite element codes. ∎
@@
Appendix B Limits of discrete Green’s functions on some degenerate meshes
In this section we revisit a class of two-dimensional, triangular meshes analyzed in Section 6 of [8], for which the ${\mathcal{P}}_{1}$-finite element solution of the Poisson equation with homogeneous Dirichlet boundary conditions satisfies the DMP, while potentially violating the angle condition in many places. The main purpose here is to expose the behavior of the discrete Green’s function as the mesh becomes degenerate. The examples in this section are easier to analyze than the one in Theorem 5.4.
The case of a single interior node.
The simplest nontrivial example of a finite element mesh is that of a triangular domain $D=\Delta ABC$ with a single vertex $P$ in the interior. We choose $P$ so that the triangle $\Delta PBC$ is isosceles with base angles equal to $\theta$, as shown in Fig. 13 (left), although this is not essential. With only one interior mesh vertex, the stiffness matrix is just the real number
${\mathbf{A}}(\theta)=a_{D}(\varphi_{P},\varphi_{P})=\int_{D}|\nabla\varphi_{p}|^{2},$
where $\varphi_{P}$ is the nodal basis function associated with $P$.
We are interested in the behavior of the discrete Green’s function as $\theta\to 0$. Cf (4.3),
${\mathbf{A}}(\theta)\geq\int_{{\mathcal{T}}_{1}}|\nabla\varphi_{p}|^{2}=\frac{\sin(\pi-2\theta)}{2\sin^{2}\theta}=\frac{2\sin\theta\cos\theta}{2\sin^{2}\theta}=\cot\theta.$
Hence,
(B.1) $\lim_{\theta\to 0}({\mathbf{A}}(\theta))^{-1}\leq\lim_{\theta\to 0}\tan\theta=0.$
This shows that the discrete Green’s function, which is represented by a number, converges to the only singular $1\times 1$ matrix, namely the number 0. Also note that the mass matrix of the mesh stays bounded as $\theta\to 0$. Therefore, given a fixed right hand side $f$ for the Poisson equation on $D$ with the mesh described above, the finite element solution $u(\theta)$ will converge to 0 as $\theta\to 0$. We remark that the limiting mesh, shown Fig. 13 (right), is a valid triangulation of $D$ with no interior nodes. Note that the Poisson problem with homogeneous Dirichlet boundary conditions can also be formulated on the associated finite element space, which contains only the function 0. Hence, the solution will be zero, which is consistent with the limit of the solutions $u(\theta)$.
Figure 13. As $\theta\to 0$ the mesh on the left converges to the mesh in the right image. When $\Delta ABC$ is an interior triangle of another triangular mesh, the partition on the right will not be a valid mesh, since the triangle below $\Delta ABC$ – not pictured – is not divided.
An embedded divided triangle.
For the second example consider a triangular mesh ${\mathcal{T}}_{h}$ on a polygonal domain $D$, and we subdivide one interior triangle as in Fig. 13 (left), which results in a refined mesh $\tilde{{\mathcal{T}}}_{h}=\tilde{{\mathcal{T}}}_{h}(\theta)$. If the discrete Green’s function is positive (or just nonnegative) on ${\mathcal{T}}_{h}$, it is shown in Section 6 of [8] that the same holds for $\tilde{{\mathcal{T}}}_{h}$, regardless of the value of $\theta>0$. For $\theta=0$, the subdivided triangle is shown in Fig. 13 (right). The resulting partition of $D$ corresponding to $\theta=0$, with $P(0)$ being the midpoint between $B$ and $C$, is not a valid triangulation, because the triangle in ${\mathcal{T}}_{h}$ below $\Delta ABC$ is not subdivided. A similar invalid partition is depicted in [4], Fig. 3.12 (right), page 80.
The question we answer here is what happens to the discrete Green’s function on $\tilde{{\mathcal{T}}}_{h}(\theta)$ as $\theta\to 0$. We will describe this limit in terms of the discrete Green’s function on ${{\mathcal{T}}}_{h}$, and we show it is represented by a rank-1 deficient matrix. We denote by ${\mathcal{B}}=\{\varphi_{1},\dots,\varphi_{n}\}$ the nodal basis associated with the interior vertices of ${\mathcal{T}}_{h}$, and let ${V}_{0}^{h}=\mathrm{span}({\mathcal{B}})$. Assume that $\varphi_{n+1}$ is the nodal basis function in $\tilde{{\mathcal{T}}}_{h}$ associated with $P=P(\theta)$ (this was called $\varphi_{P}$ from the previous example), and $\varphi_{n-1},\varphi_{n}$ correspond to $B$ and $C$, respectively. The set $\tilde{{\mathcal{B}}}={\mathcal{B}}\cup\{\varphi_{n+1}\}$ forms a hierarchical basis for the finite element space $\tilde{V}_{0}^{h}$ on $\tilde{{\mathcal{T}}}_{h}$. Cf. [8], we have $a_{D}(\varphi_{i},\varphi_{n+1})=0$ for $i=1,\dots,n$, which leads to a decoupling of the linear system representing the Poisson equation when formulated in the hierarchical basis $\tilde{{\mathcal{B}}}$. For $f\in(\tilde{V}_{0}^{h})^{*}$, the system reads:
(B.2) $\begin{bmatrix}{\mathbf{A}}&0\\ 0&\alpha\end{bmatrix}\begin{bmatrix}{\mathbf{x}}\\ x_{n+1}\end{bmatrix}=\begin{bmatrix}{\mathbf{b}}\\ b_{n+1}\end{bmatrix}$,
where $\mathbf{A}$} is the stiffness matrix of the problem in the basis $\mathcal{B}$} (hence on $V_{0}^{h}$), $\alpha=a_{D}(\varphi_{n+1},\varphi_{n+1})$, with ${{\mathbf{b}}}\in\mathbb{R}^{n}$ given by ${{\mathbf{b}}}_{i}=\left<\varphi_{i},f\right$> for $1\leq i\leq n$, and $b_{n+1}=\left<\varphi_{n+1},f\right$>. The solution of (B.2) is given by
(B.3) $u=u(\theta)=\sum_{i=1}^{n+1}x_{i}\varphi_{i},\ \ {\mathbf{x}}={\mathbf{A}}^{-1}{\mathbf{b}},\ \ x_{n+1}={b}_{n+1}/\alpha.$
If $f$ is fixed and $\theta\to 0$, it follows from (B.1) that $x_{n+1}\to 0$. Hence,
(B.4) $u_{0}=\lim_{\theta\to 0}u(\theta)=\sum_{i=1}^{n}x_{i}\varphi_{i}\in V_{h}^{0}.$
It follows that
(B.5) $u_{0}(P(0))=\frac{1}{2}(u_{0}(B)+u_{0}(C)),$
showing that the limit as $\theta\to 0$ of all the finite element solutions on $\tilde{{\mathcal{T}}}$ lie in a subspace of co-dimension 1. In other words, the degenerate partition can be considered “harmless”, as the limiting solution $u_{0}$ is simply the solution on the original mesh ${\mathcal{T}}_{h}$.
In order to describe the matrix representation $\tilde{{\mathbf{G}}}(\theta)$ of the discrete Green’s function on $\tilde{{\mathcal{T}}}$ and its limit as $\theta\to 0$, we let $f^{(i)}$ be the Dirac impulse forcing at the $i^{\mathrm{th}}$ node, for $i\leq n$. Then in (B.3) we have ${\mathbf{b}}={\mathbf{e}}_{i}$ ($i^{\mathrm{th}}$ unit vector), and $b_{n+1}=0$. Cf. (B.4) the limiting solution $u_{0}^{(i)}$ lies in $V_{h}^{0}$, and satisfies the same equation as the discrete Green’s function on ${\mathcal{T}}_{h}$. Hence, (B.5) implies that the vector representation of $u_{0}^{(i)}$ in the basis $\tilde{{\mathcal{B}}}$ is
$(\tilde{{\mathbf{g}}}^{(i)})^{T}=[({\mathbf{g}}^{(i)})^{T},\frac{1}{2}({\mathbf{g}}^{(i)}_{n-1}+{\mathbf{g}}^{(i)}_{n})$],
where ${\mathbf{g}}^{(i)}$ is the $i^{\mathrm{th}}$ column of the discrete Green’s function on ${\mathcal{T}}_{h}$.
If $f^{(n+1)}$ is the Dirac impulse forcing at $P(\theta)$, then the corresponding right-hand side in (B.3) satisfies, as $\theta\to 0$
${\mathbf{b}}=\frac{1}{2}\left({\mathbf{e}}_{n-1}+{\mathbf{e}}_{n}\right),\ \ b_{n+1}=1.$
By (B.4), its vector representation is
$(\tilde{{\mathbf{g}}}^{(n+1)})^{T}=[\frac{1}{2}({\mathbf{g}}^{(n-1)}+{\mathbf{g}}^{(n)})^{T},\frac{1}{2}({\mathbf{g}}^{(n)}_{n-1}+{\mathbf{g}}^{(n-1)}_{n})$].
Hence, if we denote by the matrix representation in $\tilde{{\mathcal{B}}}$ of the limit of the discrete function as $\theta\to 0$ is
$\tilde{{\mathbf{G}}}_{0}=\lim_{\theta\to 0}\tilde{{\mathbf{G}}}(\theta)=\begin{bmatrix}{\mathbf{G}}&\overline{g}\\ \overline{{\mathbf{g}}}^{T}&\tilde{g},\end{bmatrix}\in\mathbb{R}^{(n+1)\times(n+1)}.$
where $\mathbf{G}$} is the representation of the discrete Green’s function on the original mesh ${\mathcal{T}}_{h}$, and
$\overline{{\mathbf{g}}}=\frac{1}{2}({\mathbf{g}}^{(n-1)}+{\mathbf{g}}^{(n)}),\ \ \tilde{g}=\frac{1}{2}({\mathbf{g}}^{(n-1)}_{n-1}+{\mathbf{g}}^{(n)}_{n})$.
We used the equality ${\mathbf{g}}^{(n-1)}_{n}={\mathbf{g}}^{(n)}_{n-1}$ to compute the last diagonal entry $\tilde{g}$, which follows from the symmetry of the discrete Green’s function $\mathbf{G}$}. Hence, the last column of $\tilde{{\mathbf{G}}}_{0}$ is the average of the previous two, showing $\tilde{{\mathbf{G}}}_{0}$ has rank $n$. By contrast, the matrix ${\mathbf{T}}_{0}$ in Theorem 5.4 has rank-3 deficiency.